### v1

# Constant-Memory Strategies in Stochastic Games: Best Responses and Equilibria

Abstract
Stochastic games have become a prevalent framework for studying long-term multi-agent interactions, especially in the domain of multi-agent reinforcement learning. However, the notion of equilibria in a stochastic game is fraught with subtleties, mainly because a strategy in a stochastic game can potentially map any observed history to (possibly randomized) actions, leading to an infinitely large strategy space for each agent. In particular, it has been proven that there exists a Nash equilibrium implemented by a profile of stationary strategies. This existence result has largely drawn the community’s attention to devising multi-agent learning algorithms to reach a stationary equilibrium. Nevertheless, in real-world scenarios, particularly when one has only limited knowledge of her opponents, the assumption that all opponents will use stationary strategies is indeed too strong. To this end, in this paper, we comprehensively investigate a less restricted class of strategies called constant-memory strategies, which map constant-length history segments to actions. Despite being weaker than Turing-machine strategies, constant-memory strategies can lead to better outcomes than those under stationary strategies. We show that given a constant-memory strategy, there always exists a deterministic constant-memory strategy that uses the same length of memory as a pure strategy best response. In addition, we provide an existence result for equilibria, indicating that given any finite length of memory, there always exists a Nash equilibrium in which all agents adopt constant-memory strategies using that length of memory. However, we also show that the best response against a mixed constant-memory strategy, as if an agent is in a tournament with known opponent types, is not constant-memory and is even hard to compute. Finally, based on the aforementioned theoretical takeaways, we present simple parameterized methods to generate single-agent RL environments that can be arbitrarily difficult to learn, achievable through slight modifications of the original multi-agent environment. We hope this work significantly deepens the understanding of theoretical issues in single-agent planning under multi-agent systems and uncovers the connection between decision models in single-agent scenarios and those in multi-agent scenarios.
Keywords: Stochastic games; Bounded rationality; Best response; Restricted memory; Reinforcement learning
1 Introduction
Various real-world situations that involve long-term interactions among a group of participants can be modeled as stochastic games, such as negotiation between multiple stakeholders [8, 17, 7], bidding and mechanism design in repeated auctions [25, 9, 22, 21, 47, 43], multi-agent teamwork [42, 34, 46], and even human-robot collaboration [54, 37]. Stochastic games, also known as Markov games, model the interactions of these multi-agent systems as a Markov chain over a set of states, where the transitions are triggered by joint actions and are potentially stochastic.
The formalization of stochastic games was first proposed in Shapley’s seminal work [41]. A perfectly rational agent in a stochastic game is supposed to make use of all past histories to determine the next action, and therefore, the notion of strategies, defined as mappings from all possible histories to actions, is inherently complex. The fact that there are infinitely many strategies prohibits the direct application of Nash’s theorem for establishing any existence result of equilibria. However, the stationary transitions of stochastic games inevitably draw attention to a highly special subclass of time-independent and memoryless strategies that only consider the current states while discarding all past histories, termed stationary strategies. Indeed, the existence of equilibria formed by stationary strategies in $n$-player general-sum stochastic games was later proven by Fink [18] and Takahashi [48], under mild assumptions. Despite being highly restricted in terms of expressiveness, the notion of stationary strategies has enabled the community to practically investigate some complex real-world applications, particularly by resorting to multi-agent reinforcement learning (MARL) techniques, as advocated by Littman [29] and implemented in a line of subsequent work [30, 19, 39, 51].
Notably, one would naturally expect strategies in other less restricted forms that can encode a broader class of behavioral patterns, hoping to achieve better payoff outcomes. For example, in the Iterated Prisoner’s Dilemma (IPD), if only stationary strategies are considered, there is a unique Nash equilibrium where both players choose Defect all the time, resulting in the lowest overall payoff. However, even with the ability to remember only one past action played by the opponent, the well-known Tit-For-Tat (TFT) strategy (having Cooperate played in the first step) can be devised. One can easily see that if both players adopt the TFT strategy, they will follow a trajectory of both playing Cooperate throughout the game, resulting in a Nash equilibrium with the highest possible payoff. Apart from other forms of representation, such as strategies represented as finite automata [40, 10, 56] and even Turing machines [33, 27, 35, 13], we focus our main effort on investigating the notion of constant-memory strategies, i.e., mappings from history segments of bounded lengths to actions, mainly because it directly relates to the concept of bounded rationality [45] in general, and is highly implementable using function approximators like Recurrent Neural Networks [24, 15] and Transformers [49] in practice.
In this paper, we comprehensively study the theoretical properties associated with constant-memory strategies. We begin by presenting the following two results:
  1. 1.
    A Characterization of Best Responses: Given a constant-memory strategy profile adopted by the opponents, there always exists a deterministic constant-memory strategy that makes use of the same length of memory acting as a pure strategy best response.
  2. 2.
    An Existence Result of Equilibria: Given any finite length of memory, there always exists a Nash equilibrium where all agents adopt constant-memory (but not necessarily deterministic) strategies using that length of memory.
As a side benefit of using memories of constant length, any strategy that uses a shorter memory can always be implemented by one that uses a longer memory. Therefore, the above two results directly imply that any NE formed by shorter-length-memory strategies can be transformed into an NE formed by longer-length-memory strategies, suggesting that the longer the memory used by the strategies, the richer the equilibria one can potentially expect.
Additionally, we provide further results about best responses against mixed strategies, mathematically defined as those sampled from a set of support strategies with certain probabilities. This is associated with broad applications in the domain of opponent modeling [12, 3, 5, 55], particularly for type-based methods [3, 4, 2, 55]. However, we demonstrate that:
  1. 1.
    An Negative Result on Strategy Equivalence: An opponent with a mixed constant-memory strategy may not correspond to an equivalent opponent with a single (behavioral) constant-memory strategy in terms of achieving the same payoff.
  2. 2.
    A Negative Result on Best Responses: The best response against a mixed constant-memory strategy is not necessarily constant-memory, and computing such best responses is computationally hard.
Finally, we translate the aforementioned theoretical insights into a simple generative framework to instantiate single-agent learning environments that can be arbitrarily difficult to learn a good policy. This, in turn, reveals the potential sources of hardness underlying some single-agent decision-making problems.
2 Stochastic Games
The whole system where the agents interact is modelled as a stochastic game (SG, also known as Markov games), which can be seen as an extension of both normal-form games (to dynamic situations with stochastic transitions) and Markov decision processes (to strategic situations with multiple agents). A stochastic game is a 5-tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},T,R\rangle$ given as follows,
  1. 1.
    $\mathcal{N}$ is a finite set of $n$ agents.
  2. 2.
    $\mathcal{S}$ is a finite set of (environmental) states.
  3. 3.
    $\mathcal{A}=\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ is a set of joint actions, where $\mathcal{A}_{i}$ is the action set of agent $i$. In particular, we write $a_{i}$ as the action of agent $i$ and the one without any subscript $a=(a_{i},a_{-i})$ as the joint action.
  4. 4.
    $T:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\mathcal{A}_{n}\mapsto\Delta(\mathcal{S})$ defines stochastic transitions among states.
  5. 5.
    $R_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\mathcal{A}_{n}\mapsto\mathbb{R}$ denotes the immediate rewards for agent $i$.
To define best responses and hence equilibria, we need to first define strategies and objectives.
Assuming complete observability and perfect recall, a perfectly rational agent should utilize the entire history, while in memory-restricted cases, an agent can only devise strategies based on past memories of finite lengths. We denote the space of all possible histories of length $K\in\mathbb{N}$ as $\mathcal{H}^{K}=(\mathcal{S}\times\mathcal{A})^{K}$. In particular, when $K=0$, we have $\mathcal{H}^{0}=\emptyset$ meaning that no history can be utilized. Then, given any non-negative integer $K$, a $K$-memory strategy for agent $i$ is a mapping from all possible histories with lengths less then or equal to $K$ and the current states to (possibly randomized) actions, mathematically denoted as $\pi_{i}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$ where $\mathcal{H}^{\leq K}\triangleq\cup_{k=0}^{K}\mathcal{H}^{k}$. Let $\Pi_{i}^{K}$ denote the set of all such $K$-memory strategies for agent $i$. For convenience, we let $\mathcal{H}^{\infty}=(\mathcal{S}\times\mathcal{A})^{*}$ denote the set of complete histories that an agent with perfect recall can possibly memorize, and therefore, $\Pi_{i}^{\infty}$ is the set of all possible infinite-memory strategies for agent $i$ of the form $\pi_{i}:(\mathcal{S}\times\mathcal{A})^{*}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$. A direct consequence is that $\Pi_{i}^{K}\subseteq\Pi_{i}^{K^{\prime}}\subseteq\Pi_{i}^{\infty}$ for any non-negative $K\leq K^{\prime}$. Among them, one of the most popular class of strategies is $\Pi_{i}^{0}$, termed stationary strategies. Note that an agent capable of performing infinite-memory strategies can deliberately adopt a constant-memory strategy.
The objective for each agent is to maximize its accumulated discounted rewards (a.k.a. the discounted-payoff scenario, as opposed to the average-payoff scenario). We let $R_{i,t}$ denote the reward signaled to agent $i$ at step $t$, similarly for $S_{t}$ and $a_{i,t}$, then the overall utility under a policy profile $(\pi_{i},\pi_{-i})$ starting from any arbitrary state $S\in\mathcal{S}$ is
$u_{i}(S;\pi_{i},\pi_{-i})=\mathbb{E}_{(\pi_{i},\pi_{-i})}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R_{i,t}\Big{|}S_{0}=S\Big{]}$ (1)
$\pi_{i}$ is said to be the best response of $\pi_{-i}$, denoted as $\pi_{i}\in BR(\pi_{-i})$, if
$\forall S\in\mathcal{S},\pi_{i}^{\prime}\in\Pi_{i}^{\infty},u_{i}(S;\pi_{i},\pi_{-i})\geq u_{i}(S;\pi_{i}^{\prime},\pi_{-i})$ (2)
Note that, to compare the values of two strategy profiles, one must ensure that the limit of the right-hand side (RHS) in Equation (1) exists in the first place. Also note that, some pairs of $\pi_{i}$ and $\pi_{i}^{\prime}$ may not be comparable in the above sense, as as this comparison requires value dominance across all possible states.
3 K-Memory Best Responses and Equilibria
One should be aware of the following fact for single-agent Markov Decision Processes (MDPs) [38] in the first place, which will be considered as a lemma for the remainder of this note.
Lemma 1.
For a (single-agent) MDP $\langle S,A,T,R,\gamma\rangle$, the following two are equivalent,
  1. 1.
    Search a policy $\pi:S\mapsto\Delta(A)$ that maximizes $\mathbb{E}_{\pi}\sum_{t=0}^{\infty}[\gamma^{t}R_{t}]$, for any initial $s\in S$.
  2. 2.
    Solve the Bellman optimality equation, and then extract the policy out of the value function
    $\forall s\in S,v_{*}(s)=\max_{a\in A}\Big{[}R(s,a)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)v_{*}(s^{\prime})\Big{]}$
Assumption 1.
We assume that agents are independent of each other and rewards are bounded.
Theorem 1.
Given $\pi_{-i}\in\Pi_{-i}^{0}$, i.e. all other agents are adopting stationary strategies, it is sufficient for agent $i$ to best respond with a stationary strategy as well.
Proof.
As the state transition is stationary and the opponents are stationary, agent $i$ is therefore positioned in an induced MDP from its own perspective as follows, denoted as $\mathcal{M}^{0}(\pi_{-i})=\langle\mathcal{S},\mathcal{A}_{i},T^{0}_{\pi_{-i}},R^{0}_{\pi_{-i}},\gamma\rangle$,
  • The set of states $\mathcal{S}$, the set of control actions $\mathcal{A}_{i}$, and the discount factor $\gamma$ inherit from the original setup of the stochastic game,
  • Transition probabilities: $T^{0}_{\pi_{-i}}(S^{\prime}|S,a_{i})\triangleq\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\pi_{-i}(a_{-i}|S)$,
  • Rewards: $R^{0}_{\pi_{-i}}(S,a_{i})\triangleq\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi_{-i}(a_{-i}|S)$,
That is, the states from agent $i$’s perspective are still the environment states with the transitions compounded with the policies of the other agents acting as stationary noise. One can easily prove $T^{0}_{\pi_{-i}}$ is still a valid stochastic transition function, i.e., $\sum_{S^{\prime}}T^{0}_{\pi_{-i}}(S^{\prime}|S,a_{i})=1$ for all $a_{i}\in\mathcal{A}_{i}$.
Note that among all the optimal solutions of $\mathcal{M}^{0}(\pi_{-i})$, there is a stationary (and deterministic) policy $\pi_{opt}:\mathcal{S}\mapsto\mathcal{A}_{i}$ (cf. [38]), which corresponds to the stationary (and pure) strategy best response of agent $i$ againt $\pi_{-i}$.
We also point out that the best response can be of the form $\mathcal{S}\mapsto\Delta(\mathcal{A})$, which is called a randomized policy in control theory, or a behavior strategy (instead of a mixed strategy, as we will elaborate later in Section 4) in game theory. ∎
Similarly, we can extend this results to the case where all the opponents are equipped with constant-memory strategies, and the non-negative (and finite) memory length is the same for all opponents. We provide detailed proofs below, although the intuition behind them is quite straightforward. One can find similar proof techniques by induction in [38] (cf. Chapter 5.5).
Theorem 2.
Given $\pi_{j}\in\Pi_{j}^{K}$ for all $j\neq i$, i.e. all the other agents are adopting constant-memory strategies with the same finite memory length $K$, it is sufficient for agent $i$ to best respond with a $K$-memory strategy as well.
Proof.
Given an SG, and an opponent strategy profile $\pi_{-i}^{\infty}\in\Pi_{-i}^{\infty}$, the induced MDP in general is $\mathcal{M}^{\infty}(\pi_{-i}^{\infty})=\langle\mathcal{H}^{\infty}\times\mathcal{S},\mathcal{A}_{i},T^{\infty}_{\pi_{-i}},R^{\infty}_{\pi_{-i}},\gamma\rangle$,
  • $\mathcal{A}_{i}$ and $\gamma$ inherit from the previous setup,
  • A state is now consisting the whole history plus the current environment state, i.e. ×$\mathcal{H}^{\infty}\times\mathcal{S}$,
  • Transitions are now made also for the complete histories, as we have
    $Pr(a_{-i},S^{\prime}|H,S,a_{i})=T(S^{\prime}|S,a)\pi_{-i}^{\infty}(a_{-i}|H,S)$
    Therefore, for $(H^{\prime},S^{\prime})$, $(H,S)\in\mathcal{H}^{\infty}\times\mathcal{S}$,
    $T^{\infty}(H^{\prime},S^{\prime}|H,S,a_{i})\triangleq\begin{cases}T(S^{\prime}|S,a)\pi_{-i}^{\infty}(a_{-i}|H,S),&\text{ if }H^{\prime}=[H,S,(a_{i},a_{-i})]\\ 0,&\text{ otherwise}\\ \end{cases}$
    where $[H,S,(a_{i},a_{-i})]$ means to concatenate the existing history and the latest state-action tuple, which is a deterministic operation.
  • Rewards on the complete histories: $R^{\infty}(H,S,a_{i})\triangleq\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi_{-i}^{\infty}(a_{-i}|H,S)$,
The above $\mathcal{M}^{\infty}$ is trivially a valid MDP because transitions are made among complete state trajectories where the Markov property must hold.
Now we will show that if there exists a $\pi_{-i}^{K}\in\Pi_{-i}^{K}$, such that
$\pi_{-i}^{\infty}(a_{i}|(H^{K},H^{-}),S)=\pi_{-i}^{K}(a_{i}|H^{K},S)$ (3)
where $H^{K}=H[-\min\{K,len(H)\}:]$ and $H^{-}=H[:-\min\{K,len(H)\}]$ (the latest $K$ historical records and the remaining prefix), then for the control policy of this MDP, it is sufficient for agent $i$ to restrict the attention to $\Pi_{i}^{K}$ instead of general $\Pi_{i}^{\infty}$. More specifically, given an $\pi_{i}^{\infty}\in\Pi_{i}^{\infty}$, it is possible to construct a memory-restricted alternative $\pi_{i}^{K}$ such that the following target equation holds
$Pr^{\pi_{i}^{K}}(H^{K},S,a_{i})=Pr^{\pi_{i}^{\infty}}(H^{K},S,a_{i})$ (4)
where $Pr^{\pi}$ means the probability under the particular policy $\pi$. The above proof target is sufficient in terms of seeking for an equivalent solution because it directly pertains to the reward function. We will show that such a strategy for agent $i$ can be constructed by the following, i.e. by marginalizing over histories happened earlier then $K$ steps ago,
$\pi_{i}^{K}(a_{i}|H^{K},S)=\sum_{H^{-}}\pi_{i}^{\infty}(a_{i}|H^{K},H^{-},S)Pr(H^{-})$ (5)
We will prove this equation by induction.
For the base case, when $|H|=0$ which simply means $S$ is the initial state, then Equation (4) obviously holds.
For the inductive case, we hypothesize that the following holds for all possible $(\hat{H},\hat{S})$ with $|\hat{H}|=t-1$,
$Pr^{\pi_{i}^{K}}(\hat{H}^{K},\hat{S},a_{i})=Pr^{\pi_{i}^{\infty}}(\hat{H}^{K},\hat{S},a_{i})$
Because of Equation (3), we have
$\begin{split}Pr(a_{-i},S^{\prime}|H,S,a_{i})&=T(S^{\prime}|S,a)\pi_{-i}^{\infty}(a_{-i}|(H^{K},H^{-}),S)\\ &=T(S^{\prime}|S,a)\pi_{-i}^{K}(a_{-i}|H^{K},S)\\ &=Pr(a_{-i},S^{\prime}|H^{K},S,a_{i})\end{split}$ (6)
Then for $|H|=t$, we have
$\begin{split}Pr^{\pi_{i}^{K}}(H^{K},S)&=\sum_{(\hat{H}^{K},\hat{S})}\sum_{a_{i}^{\prime}}Pr^{\pi_{i}^{K}}(\hat{H}^{K},\hat{S},a_{i}^{\prime}){\color[rgb]{0,.5,.5}Pr^{\pi_{i}^{K}}(H^{K},S|\hat{H}^{K},\hat{S},a_{i}^{\prime})}\\ &=\sum_{(\hat{H}^{K},\hat{S})}\sum_{a_{i}^{\prime}}Pr^{\pi_{i}^{\infty}}(\hat{H}^{K},\hat{S},a_{i}^{\prime}){\color[rgb]{0,.5,.5}Pr^{\pi_{i}^{\infty}}(H^{K},S|\hat{H}^{K},\hat{S},a_{i}^{\prime})}\\ &=Pr^{\pi_{i}^{\infty}}(H^{K},S)\end{split}$ (7)
The second equality directly follows from the inductive hypothesis and Equation (6). Note that, for the terms in teal, it does not matter which rollout policy is used, as $a_{i}^{\prime}$ is conditioned.
Finally, we have
$\begin{split}Pr^{\pi_{i}^{K}}(H^{K},S,a_{i})&=Pr^{\pi_{i}^{K}}(H^{K},S)\times Pr^{\pi_{i}^{K}}(a_{i}|H^{K},S)\\ &=Pr^{\pi_{i}^{K}}(H^{K},S)\times\pi_{i}^{K}(a_{i}|H^{K},S)\\ &=Pr^{\pi_{i}^{K}}(H^{K},S)\times\sum_{H^{-}}\pi_{i}^{\infty}(a_{i}|H^{K},H^{-},S)Pr(H^{-})\\ &=Pr^{\pi_{i}^{\infty}}(H^{K},S)\times Pr^{\pi_{i}^{\infty}}(a_{i}|H^{K},S)\\ &=Pr^{\pi_{i}^{\infty}}(H^{K},S,a_{i})\end{split}$
The third equality holds according to Equation (5), and the fourth equality directly follows from Equation (7).
As a conclusion, and perhaps from a more direct angle, one can see that given $\pi_{j}\in\Pi_{j}^{K}$ for all $j\neq i$, agent $i$ is then faced with an MDP with environment states augmented by finite-length histories, denoted as $\mathcal{M}^{K}(\pi_{-i})=\langle\mathcal{H}^{\leq K}\times\mathcal{S},\mathcal{A}_{i},T^{K}_{\pi_{-i}},R^{K}_{\pi_{-i}},\gamma\rangle$,
  • $\mathcal{A}_{i}$ and $\gamma$ inherit from the previous setup,
  • A state is now consisting the past $K$ steps plus the current environment state, i.e. $\mathcal{H}^{\leq K}\times\mathcal{S}$,
  • Transitions are now made for the augmented states, i.e., for $(H^{\prime},S^{\prime})$, $(H,S)\in\mathcal{H}^{\leq K}\times\mathcal{S}$,
    $T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})\triangleq\begin{cases}T(S^{\prime}|S,a)\pi_{-i}(a_{-i}|H,S),&\text{ if }H^{\prime}=slide_{K}(H,S,(a_{i},a_{-i}))\\ 0,&\text{ otherwise}\\ \end{cases}$
    where $slide_{K}(H,S,(a_{i},a_{-i}))$ means to discard the earliest step if it is more then $K$ steps away, and append the latest state and action profile.
  • Rewards for each in $\mathcal{H}^{\leq K}\times\mathcal{S}$: $R^{K}_{\pi_{-i}}(H,S,a_{i})\triangleq\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi_{-i}(a_{-i}|H,S)$,
Again, among all the optimal solutions of $\mathcal{M}^{K}(\pi_{-i})$, there is a stationary (and deterministic) policy, i.e. $\pi_{opt}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\mathcal{A}_{i}$, which corresponds to the $K$-memory (and pure) strategy best response of agent $i$ against $\pi_{i}\in\Pi_{-i}^{K}$. ∎
A natural next question is: what if the opponents are heterogeneous in their adoption of constant-memory strategies with different memory lengths? The following summarizes a related corollary.
Theorem 3.
Given $\pi_{j}\in\Pi_{j}^{K_{j}}$ for all $j\neq i$, i.e. all the other agents are adopting constant-memory strategies but with varying memory lengths, it is sufficient for agent $i$ to best respond with a $(\max\{K_{j}\}_{j\neq i})$-memory strategy.
Proof sketch.
We omit the rigorous proof here. The only difference from the previous ones is that when agent $i$ is looking ahead for possible successor states, it should take into account the opponent that is equipped with the longest memory, since the history may appear distinguishable only to that particular agent while remaining indistinguishable to the rest. That is, the transition from agent $i$’s perspective is primarily controlled by the opponent with the longest memory.
Therefore, the states of the induced MDP will be over the space of $\mathcal{H}^{\leq\max\{K_{j}\}_{j\neq i}}\times\mathcal{S}$, and then there will be an optimal solution as a stationary (and deterministic) mapping $\mathcal{H}^{\leq\max\{K_{j}\}_{j\neq i}}\times\mathcal{S}\mapsto\mathcal{A}_{i}$, which corresponds to the $(\max\{K_{j}\}_{j\neq i})$-memory (and pure) strategy best response of agent $i$. ∎
As best responses are well established, we will examine whether an equilibrium exists when everyone is best responding to one another.
Definition 1 (Nash Equilibrium).
A strategy profile $\{\pi_{i}^{*}\}_{i\in\mathcal{N}}$ is a Nash equilibrium (NE) if
$\forall i\in\mathcal{N},\pi_{i}^{*}\in BR(\pi_{-i}^{*})$
We first need the following lemma. It is important to note that the following lemma only asserts the existence of a fixed point, but does not guarantee the presence of a contraction mapping.
Lemma 2 (Brouwer’s fixed-point theorem).
Let $\Delta=\prod_{l=1}^{L}\Delta_{m_{l}}$ where each $\Delta_{m_{l}}$ is a simplex in $\mathbb{R}^{m_{l}+1}$. Let $f:\Delta\mapsto\Delta$ be continuous. Then $f$ has a fixed point.
Theorem 4.
There exists a Nash equilibrium when the agents are all adopting $K$-memory strategies, for any arbitrary non-negative finite $K$.
Proof.
Given a non-negative finite $K$, to establish a Nash equilibrium we need to prove there is a solution to the system of equations defined by
$\forall i\in\mathcal{N},\pi_{i}\in\Pi_{i}^{K}\land\pi_{i}\in BR(\pi_{-i})$
More specifically, the following $n$ Bellman optimality equations should be satisfied simultaneously for any $(H,S)\in\mathcal{H}^{\leq K}\times\mathcal{S}$,
$\begin{split}v_{1}(H,S)&=\max_{a_{1}\in\mathcal{A}_{1}}\Big{[}R^{K}_{\pi_{-1}}(H,S,a_{1})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-1}}(H^{\prime},S^{\prime}|H,S,a_{1})v_{1}(H^{\prime},S^{\prime})\Big{]}\\ &\vdots\\ v_{n}(H,S)&=\max_{a_{n}\in\mathcal{A}_{n}}\Big{[}R^{K}_{\pi_{-n}}(H,S,a_{n})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-n}}(H^{\prime},S^{\prime}|H,S,a_{n})v_{n}(H^{\prime},S^{\prime})\Big{]}\\ \end{split}$ (8)
We prove the entire theorem by two steps:
  1. 1.
    We first show that, given any strategy profile $\{\pi_{i}\}_{i\in\mathcal{N}}$, a unique solution, i.e., a set of values $\{v_{i}(\cdot,\cdot)\}_{i\in\mathcal{N}}$, for Equation (8) is guaranteed to exist.
  2. 2.
    We then show that there exists such a strategy profile $\{\pi_{i}^{*}\}_{i\in\mathcal{N}}$ such that the values of $\pi_{i}^{*}$ satisfy the $i$-th equation in Equation (8) simultaneously for all $i\in\mathcal{N}$.
Step 1.
Let $\mathcal{V}$ denote the vector space of all possible value functions, where each $v\in\mathcal{V}$ is a function $\mathcal{N}\times\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\mathbb{R}$ (slightly reloading the notation $v_{i}(H,S)$). Let $\Xi:\mathcal{V}\mapsto\mathcal{V}$ denote the (multi-agent) Bellman optimality operator given as follows,
$\Xi(v)(i,H,S)=\max_{a_{i}\in\mathcal{A}_{i}}\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})v(i,H^{\prime},S^{\prime})\Big{]}$
For the rest, we write $\Xi_{v}$ interchangeably with $\Xi(v)$ for better presentation. We use the infinity norm as the distance measure, defined as $\|v\|_{\infty}=\max_{x}|v(x)|$ for $v\in\mathcal{V}$. We then show for any two vectors $u,v\in\mathcal{V}$, we have $\|\Xi(u)-\Xi(v)\|_{\infty}\leq\gamma\|u-v\|_{\infty}$.
For $(i,H,S)$ such that $\Xi_{u}(i,H,S)\geq\Xi_{v}(i,H,S)$, we choose
$a_{i}^{*}\in\arg\max_{a_{i}\in\mathcal{A}_{i}}\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})u(i,H^{\prime},S^{\prime})\Big{]}$
then,
$\begin{split}|\Xi_{u}(i,H,S)-\Xi_{v}(i,H,S)|&=\Xi_{u}(i,H,S)-\Xi_{v}(i,H,S)\\ &=\max_{a_{i}\in\mathcal{A}_{i}}\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})u(i,H^{\prime},S^{\prime})\Big{]}\\ &-\max_{a_{i}\in\mathcal{A}_{i}}\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})v(i,H^{\prime},S^{\prime})\Big{]}\\ &\leq\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i}^{*})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i}^{*})u(i,H^{\prime},S^{\prime})\Big{]}\\ &-\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i}^{*})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i}^{*})v(i,H^{\prime},S^{\prime})\Big{]}\\ &=\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i}^{*})[u(i,H^{\prime},S^{\prime})-v(i,H^{\prime},S^{\prime})]\\ &\leq\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i}^{*})|u(i,H^{\prime},S^{\prime})-v(i,H^{\prime},S^{\prime})|\\ &\leq\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i}^{*})\|u-v\|_{\infty}=\gamma\|u-v\|_{\infty}\end{split}$
One can use a similar technique to also conclude $|\Xi_{u}(i,H,S)-\Xi_{v}(i,H,S)|\leq\gamma\|u-v\|_{\infty}$ for the case when $\Xi_{u}(i,H,S)<\Xi_{v}(i,H,S)$. Overall, we have
$\|\Xi(u)-\Xi(v)\|_{\infty}=\max_{i,H,S}|\Xi_{u}(i,H,S)-\Xi_{v}(i,H,S)|\leq\gamma\|u-v\|_{\infty}$
Thus, $\Xi$ is a contraction mapping, and it naturally follows that $\Xi$ has only one unique fixed point.
Step 2.
To prove that an equilibrium point exists, we will first construct a mapping to iteratively refine the strategies, and then show that there is a bijection between the fixed points of this mapping and the solutions to the aforementioned system of equations.
From each agent $i$’s perspective, with the opponent’s strategies given as $\pi_{-i}$, she shall evaluate the value of her own strategy by the Bellman expectation equation, i.e.,
$v_{i}|_{\pi_{i}}^{\pi_{-i}}(H,S)=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|H,S)\Big{[}R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})v_{i}|_{\pi_{i}}^{\pi_{-i}}(H^{\prime},S^{\prime})\Big{]}$
where $v_{i}|_{\pi_{i}}^{\pi_{-i}}$ is the value function evaluated using $\pi_{i}$ under $\pi_{-i}$. We here omit the proof that this equation has a unique solution serving as the evaluated values. We define the refinement as
$\phi_{i,a_{i}}(\pi_{i},H,S)=\max\{0,Q_{i}|_{\pi_{i}}^{\pi_{-i}}(H,S,a_{i})-v_{i}|_{\pi_{i}}^{\pi_{-i}}(H,S)\}$ (9)
where $Q_{i}|_{\pi_{i}}^{\pi_{-i}}(H,S,a_{i})=R^{K}_{\pi_{-i}}(H,S,a_{i})+\gamma\sum_{H^{\prime},S^{\prime}}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})v_{i}|_{\pi_{i}}^{\pi_{-i}}(H^{\prime},S^{\prime})$.
A (refinement) mapping $\Gamma:\{\Pi_{i}^{K}\}_{i\in\mathcal{N}}\mapsto\{\Pi_{i}^{K}\}_{i\in\mathcal{N}}$ is constructed as
for $\text{for each }i\in\mathcal{N},\ \pi_{i}(a_{i}|H,S)\mapsto\frac{\pi_{i}(a_{i}|H,S)+\phi_{i,a_{i}}(\pi_{i},H,S)}{\sum_{b_{i}\in\mathcal{A}_{i}}\pi_{i}(b_{i}|H,S)+\phi_{i,b_{i}}(\pi_{i},H,S)}$ (10)
By Lemma 2, $\Gamma$ has at least one fix point, as each state-action mapping is a simplex $\Delta_{|\mathcal{A}_{i}|-1}$.
If $\{\pi_{i}\}_{i\in\mathcal{N}}$ is already an NE, then all $\phi$’s will be zeros, making it a fixed point of $\Gamma$.
Conversely, we can show that any arbitrary fixed point of $\Gamma$, say $\{\hat{\pi}_{i}\}_{i\in\mathcal{N}}$, is also an NE. As $v$-functions are averaging over $Q$-functions, there must exist an $a_{i}^{\prime}\in\mathcal{A}_{i}$, such that (fixing an $(H,S)$)
$\hat{\pi}_{i}(a_{i}^{\prime}|H,S)>0,\text{ and }Q_{i}|_{\hat{\pi}_{i}}^{\hat{\pi}_{-i}}(H,S,a_{i}^{\prime})-v_{i}|_{\hat{\pi}_{i}}^{\hat{\pi}_{-i}}(H,S)\leq 0$
By Equation (9), we have $\phi_{i,a_{i}^{\prime}}(\hat{\pi}_{i},H,S)=0$. Given that $\{\hat{\pi}_{i}\}_{i\in\mathcal{N}}$ is already a fixed point, by definition $\{\hat{\pi}_{i}\}_{i\in\mathcal{N}}=\Gamma(\{\hat{\pi}_{i}\}_{i\in\mathcal{N}})$, and therefore, the normalization term (the denominator) must be exactly one. Due to the fact that $\phi$’s are always non-negative, then we can conclude that for all $b_{i}\in\mathcal{A}_{i}$, it must be the case $\phi_{i,b_{i}}(\hat{\pi}_{i},H,S)=0$. Hence, $v_{i}|_{\hat{\pi}_{i}}^{\hat{\pi}_{-i}}(H,S)\geq\max_{a_{i}\in\mathcal{A}_{i}}{Q_{i}|_{\hat{\pi}_{i}}^{\hat{\pi}_{-i}}(H,S,a_{i}^{\prime})}$. Consequently, the equality shall hold. One can then see it it exactly the case when the aforementioned system of Bellman optimality equations (Equation (8)) is satisfied. ∎
The above existence result indicates the following. Consider two agents playing a stochastic game, where agent 1 employs a two-memory strategy and agent 2 uses a three-memory strategy. If agent 1 asserts that she will adhere to her two-memory strategy, agent 2 may identify another two-memory strategy as a best response, potentially yielding the same payoff but allowing for memory saving. Conversely, if agent 2 can convince agent 1 that she will maintain her three-memory strategy, agent 1 may find it advantageous to switch to a three-memory strategy as a better response.
Note that the above theorem is not a consequence of directly invoking Nash’s existence theorem, as we are not discussing mixed strategies, a topic that will be addressed in the following section.
Another benefit of constant-memory strategies is that any $K$-memory strategy can be implemented using a $K^{\prime}$-memory strategy, provided that $K^{\prime}\geq K$, by simply utilizing the most recent $K$ historical records. Thus, we arrive at the following corollary.
Corollary 1.
Any payoff profile that is reached by an NE under a $K$-memory strategy profile can also be realized by another NE under a $K^{\prime}$-memory strategy profile, as long as $K^{\prime}\geq K$.
4 Best Responses to Mixed Strategies: A Tournament Perspective
We have answered the question that best responding to a single (possibly randomized) constant-memory strategy will lead to another constant-memory strategy. The next natural question is: what is the best response against a set of constant-memory strategies played over a pre-specified distribution (i.e. mixed strategies)? An additional related question is: can a mixed strategy be transformed into a singleton constant-memory strategy? If this is possible, then the best response must also be a constant-memory strategy.
In the following two subsections, we will show:
  1. 1.
    In repeated games, if an agent faces an opponent using a mixed zero-memory strategy, then it will yield the same expected utility for her to play against an opponent with a transformed singleton zero-memory strategy.
  2. 2.
    In general, when the game involves multiple states or the opponent employs a non-zero-memory strategy, then the best response will be hard to compute and time-dependent, which can not be encoded within a finite-memory strategy. This, in turn, implies that the opponent’s mixed strategy cannot be equivalently transformed into a singleton constant-memory strategy.
4.1 Mixed Strategies Are Something More Than Behavioral Strategies
We first emphasize the notion of match. When we say an agent $i$ adopts a $K$-memory strategy, it means agent $i$ will select one strategy $\pi_{i}\in\Pi_{i}^{K}$ right before a match begins. Once she has “confirmed” its strategy, she will not deviate to any other ones during the match until the termination. Note that some strategies, especially ones in $\Pi_{i}^{\infty}$, may be semantically interpreted as “learning” or “evolving” strategies, as they gradually alter the decisions based on the accumulated observations, but each of them is still a singleton strategy out of the strategy space $\Pi_{i}^{\infty}$. From the perspective of a single agent, we may refer to it as an episode, interchangeably with the term match, as is commonly done in the context of MDPs. The overall utility will be calculated as the expectation over all possible matches.
Now we are ready to explain the difference between a behavioral strategy and a mixed strategy. A behavioral strategy is a randomized policy (for agent $i$ and of $K$-memory) $\pi_{i}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$, given that $K\in\mathbb{N}$. By definition, a pure strategy is also a behavioral strategy. A mixed strategy (for agent $i$ and of $K$-memory) first specifies its support $\Pi^{K+}_{i}\subseteq\Pi_{i}^{K}$, where each $\pi_{i}^{\iota}\in\Pi^{K+}_{i}$ will be selected with a positive probability $p_{\iota}$, before every match begins. Therefore, we use a tuple $(\Pi^{K+}_{i},\vec{p})$ to denote a mixed strategy for agent $i$. Intuitively, when an agent is playing against a mixed strategy $(\Pi^{K+}_{j},\vec{p})$, it simply means this particular agent will encounter an opponent using the behavioral strategy (or, of type) $\pi_{j}^{\iota}\in\Pi^{K+}_{j}$ for a fraction $p_{\iota}$ of the whole time.
One may be particularly interested in a special form of strategies, which is the behavioral strategy obtained by state-wise randomization over the probability distribution provided by the mixed strategy.
Definition 2 (Mixed-strategy-induced behavior strategy).
Given a mixed strategy $(\Pi^{K+}_{i},\vec{p})$, we define $\omega_{(\Pi^{K+}_{i},\vec{p})}$ as the behavior strategy induced by the mixed strategy $(\Pi^{K+}_{i},\vec{p})$,
for $\text{for each }(H,S)\in\mathcal{H}^{\leq K}\times\mathcal{S},\ \omega_{(\Pi^{K+}_{i},\vec{p})}(a_{i}|H,S)=\sum_{\iota}p_{\iota}\pi_{i}^{\iota}(a_{i}|H,S)$
The intuition underneath is that, instead of randomly choosing one of the support strategies at the beginning and adhering to it, we also allow an agent to switch to another one up to the same probability at each timestep during the play, which leads to a singleton strategy that randomizes over each support strategy at every state. One can see that if the original strategy is a mixed one over a set of $K$-memory support strategies, then the induced behavioral strategy, according to Definition 2, will still be a $K$-memory strategy, and its best response will also be a $K$-memory strategy, as stated in Theorem 2.
We first show that in a special case where the stochastic game is degraded to a repeated game and agents therein use stationary (i.e., zero-memory) strategies, a mixed strategy has the same effect as its induced behavioral strategy. However, in general, if the game involves transitions over multiple states or the opponents adopt non-zero-memory strategies, such equivalence does not necessarily apply.
Theorem 5 (Utility equivalence for repeated games).
If the stochastic game is merely a repeated game, i.e. $\mathcal{S}$ is a singleton, then an agent $i$’s overall utility when she plays against the mixed strategy $(\Pi^{0+}_{-i},\vec{p})$ will be the same as that when she plays against the induced behavioral strategy $\omega_{(\Pi^{0+}_{-i},\vec{p})}$.
Proof.
Assume agent $i$ is performing any arbitrary policy $\pi_{i}$. To compute her expected return against the mixed strategy $(\Pi^{0+}_{-i},\vec{p})$, one needs to establish the Bellman expectation equation for each MDP $\mathcal{M}(\pi^{\iota}_{-i})$ induced by the opponent strategy $\pi_{-i}^{\iota}\in\Pi^{0+}_{-i}$,
$\begin{split}V_{\iota}&=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\cdot[R_{\pi_{-i}^{\iota}}(a_{i})+\gamma V_{\iota}]\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\cdot[\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}^{\iota}(a_{-i})\cdot R_{i}(a_{i},a_{-i})+\gamma V_{\iota}]\end{split}$
where $V_{\iota}$, a shorthand for $V_{\mathcal{M}(\pi^{\iota}_{-i})}$ as $\mathcal{M}(\pi^{\iota}_{-i})$ is no longer a multi-state MDP, denotes the expected return for agent $i$ when she is playing $\pi_{i}$ against $\pi_{-i}^{\iota}$. Then, one can get the following
$V_{\iota}=\frac{\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}^{\iota}(a_{-i})R_{i}(a_{i},a_{-i})}{1-\gamma}$
The overall expected utility against this mixed strategy is therefore
$V_{mix}=\sum_{\iota}p_{\iota}\frac{\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}^{\iota}(a_{-i})R_{i}(a_{i},a_{-i})}{1-\gamma}$ (11)
Now assume she is playing against the mixed-strategy-induced behavioral strategy $\omega_{(\Pi^{0+}_{-i},\vec{p})}$, the corresponding Bellman equation is
$\begin{split}V_{beh}&=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\cdot[R_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}(a_{i})+\gamma V_{beh}]\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\cdot[\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(a_{i},a_{-i})\cdot\omega_{(\Pi^{0+}_{-i},\vec{p})}(a_{-i})+\gamma V_{beh}]\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\cdot[\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(a_{i},a_{-i})\cdot(\sum_{\iota}p_{\iota}\cdot\pi^{\iota}_{-i}(a_{-i}))+\gamma V_{beh}]\\ \end{split}$
Hence, solve the equation to get
$\begin{split}V_{beh}=\frac{\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(a_{i},a_{-i})\sum_{\iota}p_{\iota}\cdot\pi^{\iota}_{-i}(a_{-i})}{1-\gamma}\end{split}$ (12)
Comparing Equation (11) and (12), one can easily see $V_{mix}=V_{beh}$, up to different orders of summation . ∎
Theorem 6 (Utility equivalence does not hold for general stochastic games).
In general, when a stochastic game involves multiple states, an agent $i$’s overall utility when she plays against the mixed strategy $(\Pi^{0+}_{-i},\vec{p})$ is not necessarily the same as that when she plays against the induced behavioral strategy $\omega_{(\Pi^{0+}_{-i},\vec{p})}$.
Proof.
Similarly as before, to evaluate an arbitrary $\pi_{i}$ under $\mathcal{M}(\pi^{\iota}_{-i})$ , one can establish the following
$\begin{split}V_{mix}(S)&=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot Q_{\mathcal{M}(\pi^{\iota}_{-i})}(S,a_{i})\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot\Big{[}R_{\pi^{\iota}_{-i}}(S,a_{i})+\gamma\sum_{S^{\prime}}T_{\pi^{\iota}_{-i}}(S^{\prime}|S,a_{i})V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\sum_{\iota}p_{\iota}\cdot\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi^{\iota}_{-i}(a_{-i}|S)+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\pi^{\iota}_{-i}(a_{-i}|S)V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)V_{\mathcal{M}(\pi^{\iota}_{-i})}(S^{\prime})\Big{]}\\ &{\color[rgb]{0,.5,.5}\ \neq\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}+\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}V_{mix}(S^{\prime})\Big{]}}\\ \end{split}$
The last equation does not necessarily hold as one cannot simply replace $V_{\mathcal{M}(\pi^{\iota}_{-i})}$ with $V_{mix}$, as it will require to solve another totally different equation. However, this particular equation is by definition the one that $V_{beh}$ should satisfy, i.e.,
$\begin{split}V_{beh}(S)=\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i}|S)\Big{[}&\sum_{a_{i}\in\mathcal{A}_{-i}}R_{i}(S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}\\ +&\gamma\sum_{S^{\prime}}\sum_{a_{-i}\in\mathcal{A}_{-i}}T(S^{\prime}|S,a)\underbrace{\sum_{\iota}p_{\iota}\pi^{\iota}_{-i}(a_{-i}|S)}_{\omega_{(\Pi^{0+}_{-i},\vec{p})}}V_{beh}(S^{\prime})\Big{]}\end{split}$ (13)
Hence, it is not necessarily the case that $V_{mix}=V_{beh}$. ∎
Remark 1 (Singleton state but positive-length memory).
As both increasing the length of memory and the number of environment states will lead to a multi-state MDP from each individual agent’s perspective, a natural corollary is that utility equivalence between a mixed strategy and its induced behavioral strategy does not necessarily hold for $K$-memory strategies once $K$ is positive, even under repeated games.
One may further wonder whether a group of agents can form some equilibrium if all of them play mixed strategies, i.e., $\{(\Pi^{K+}_{i},\vec{p}_{i})\}_{i\in\mathcal{N}}$. The story is that, if the support $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$ and a distribution over initial states $d_{0}\in\Delta(\mathcal{S})$ can be specified in the first place, the stochastic game can be further reduced to a normal-form game $\langle\mathcal{N},\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}},\{u_{i}\}_{i\in\mathcal{N}}\rangle$ (with $u_{i}$ slighted overloaded) as follows,
  1. 1.
    The game contains all agents $\mathcal{N}$,
  2. 2.
    The action set of agent $i$ is $\Pi^{K+}_{i}$, i.e. to select a behavioral strategy therein,
  3. 3.
    The payoff of agent $i$ is
    $u_{i}(\pi_{i},\pi_{-i})=\sum_{S\in\mathcal{S}}d_{0}(S)\cdot\mathbb{E}_{(\pi_{i},\pi_{-i})}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}R_{i,t}\Big{|}S_{0}=S\Big{]}$
Under this sense and provided that the reduced game is finite, invoking Nash’s well-known existence theorem, we can conclude that there must exists a mixed strategy NE $\{\vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$. That is, given fixed supports $\{\Pi^{K+}_{i}\}_{i\in\mathcal{N}}$, no one will be strictly better off by unitarily deviating from $\{\vec{p_{i}}^{*}\}_{i\in\mathcal{N}}$ to another distribution for mixing over its support strategies. However, the application of this result remains an open (and perhaps even unjustified) problem. One idea might be promising: as we will see later, finding a behavioral strategy best response to a mixed strategy is computationally hard, but will it helps if it allows for finding a mixed strategy best response instead?
4.2 Computing Best Responses to Mixed Strategies is Intractable
We will first show that computing the best response against a mixed $K$-memory strategy reduces to optimally solving an infinite-horizon partially observable MDPs (POMDPs), so that the complexity/computability results [31, 32] can naturally follow. It turns out the reduced ones belong to a subclass of generic POMDPs, namely Contextual MDP (CMDPs), although it may not potentially imply less challenging computation. To show that such reduction does not complicate the original problem, we also construct a reduction from the computational problem of solving CMDPs back to that of computing best responses (against mixed strategies).
Theorem 7.
Given a mixed strategy profile $(\Pi^{K+}_{-i},\vec{p})$ of the opponents, computing the best response for agent $i$ can be reduced to optimally solving an infinite-horizon POMDP.
Proof.
We first construct the reduction to the corresponding POMDP. The POMDP is given as a tuple $\langle\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i},\mathcal{A}_{i},\mathcal{H}^{K}\times\mathcal{S},\mathbf{T},\mathbf{O},\mathbf{R},\gamma\rangle$,
  1. 1.
    States: $\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i}$ denote the set of underlying states. That is, a state in this POMDP is the history segment and environment state of the completely observable stochastic game augmented by the unobservable opponent strategies.
  2. 2.
    As previously, $\mathcal{A}_{i}$ is the set of available control actions of agent $i$, and $\gamma$ the discount factor.
  3. 3.
    Observations: $\mathcal{H}^{K}\times\mathcal{S}$ denote the set of observations that can be made by agent $i$.
  4. 4.
    $\mathbf{T}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\times\mathcal{A}_{i}\mapsto\Delta(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})$ denote transition function, mathematically defined as
    $\mathbf{T}\Big{(}(H^{\prime},S^{\prime},\pi_{-i}^{\prime})\Big{|}(H,S,\pi_{-i}),a_{i}\Big{)}=\begin{cases}T^{K}_{\pi_{-i}}(H^{\prime},S^{\prime}|H,S,a_{i})&\text{, if }\pi_{-i}^{\prime}=\pi_{-i}\\ 0&\text{, otherwise}\end{cases}$
  5. 5.
    $\mathbf{O}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\mapsto\mathcal{H}^{K}\times\mathcal{S}$ denote the deterministic observation function, mathematically defined as
    $\mathbf{O}\Big{(}(H,S,\pi_{-i}^{\prime})\Big{)}=(H,S)$
  6. 6.
    $\mathbf{R}:(\mathcal{H}^{K}\times\mathcal{S}\times\Pi^{K+}_{-i})\times\mathcal{A}_{i}\mapsto\mathbbm{R}$, mathematically defined as
    $\mathbf{R}\Big{(}(H,S,\pi_{-i}^{\prime}),a_{i}\Big{)}=R^{K}_{\pi_{-i}}(H,S,a_{i})$
Then, we need to show that such a reduction is correct, i.e., a solution maximizes agent $i$’ expected payoff under the stochastic game w.r.t. the opponents’ mixed strategy iff it maximizes the expected return in this reduced POMDP. The argument is made by three steps:
  1. 1.
    Given any initial state $S\in\mathcal{S}$, and any sequence of joint actions, the amount of historic information that an agent with perfect recall can possibly obtain will be the same at each timestep under both models.
    • The accumulated information that agent $i$ in the stochastic game can gather is the following set
      $\{\vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}$
      and that in the reduced POMDP is all the historic observations
      $\begin{split}&\{\vec{q},S_{0}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),S_{1}\}\\ \cup&\{(S_{0},a_{i,0},a_{-i,0}),(S_{1},a_{i,1},a_{-i,1}),S_{2}\}\\ &\cdots\\ \cup&\{(S_{t-K},a_{i,t-K},a_{-i,t-K}),\cdots,(S_{t-1},a_{i,t-1},a_{-i,t-1}),S_{t}\}\\ =&\{\vec{q},S_{0},a_{i,0},a_{-i,0},S_{1},a_{i,1},a_{-i,1},\cdots,S_{t}\}\end{split}$
  2. 2.
    Given any initial state $S\in\mathcal{S}$, and any sequence of agent $i$’ actions, the probability of reaching the same trajectory will be the same.
    • Because the opponents’ actions are merely sampled from a constant-memory strategy.
  3. 3.
    Given any initial state $S\in\mathcal{S}$, and policy that maps from all possible historic information to actions will result in the same payoff under both models.
    • Note that in an episode (or a match), the opponents will not switch to another strategy profile, therefore, the total return/payoff will solely depend on the probabilities of each possible trajectory under the two models, which is ensured to be the same by the aforementioned two points.
One can see that the reduced POMDPs in the above theorem actually belong to a subclass of generic POMDPs, where a state consists of variables that can be directly observed and other hidden ones. Such a subclass is specially termed as Mixed observability MDPs (MOMPDs) [36, 6, 28]. It has been shown that planning algorithms originally developed for POMDPs are significantly faster for those factorized models like MOMDPs in practice. A more restricted model is the Contextual MDP (CMDP) [23, 11], which can be viewed as a special case of MOMDP where there is no transition between different hidden state variables. Although one can see that CMDP $\subseteq$ MOMDP $\subseteq$ POMDP in terms of computational hardness, the complexity results in general for the former two remain open. So far, the common conjecture is that neither CMDP nor MOMDP is significantly easier to solve than POMDP, and it is proven that optimally solving infinite-horizon POMDPs is undecidable [32].
This result highly pertains to some discussions on type-based methods for single-agent planning under the presence of multiple other agents [3, 4, 2, 55]. Albrecht and Ramamoorthy [3] characterized the general problem from a conceptual perspective, where each opponent’s strategy acts as an oracle that can be queried, but left the concrete issues regarding implementation unresolved. As a supplementary, Zhu and Lin [55] provided a spectrum of implementable planners for the stationary base case, i.e., each support strategy of the opponent’s mixed strategy is a stationary one. Here, Theorem 7 further generalizes it to constant-memory strategies, allowing all the formulations in [55] to be carried over to the entire family of constant-memory strategies..
Theorem 8.
Solving CMDP reduces to computing the best response for agent $i$ to a mixed stationary strategy profile $(\Pi^{0+}_{-i},\vec{p})$ of the opponents.
Proof.
We prove the above theorem by constructing a reduction for every CMDP instance to a BestResponse instance. Given a CMDP $\langle\mathcal{C},\mathcal{S},\mathcal{A},f_{T},f_{R},\gamma\rangle$, where
  1. 1.
    $\mathcal{C}$ is a finite set of unobservable contexts, one of which will be selected at the beginning of an episode.
  2. 2.
    $f_{T}$ and $f_{R}$ take in a context $c\in\mathcal{C}$ and output a transition function $T^{c}:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ as well as a reward function $R^{c}:\mathcal{S}\times{A}\mapsto\mathbbm{R}$, respectively. $\langle\mathcal{S},\mathcal{A},T^{c},R^{c},\gamma\rangle$ will then constitutes an MDP.
one can construct a stochastic game with two players $\langle\{1,2\},\mathcal{S},\{\mathcal{A},\mathcal{A^{\prime}}\},T_{0},\{R_{1},R_{2}\}\rangle$, where agent 1 with action set $\mathcal{A}$ is playing against agent 2 (as a context controller/switcher) with a set of stationary strategies $\{\pi_{c}:\mathcal{S}\mapsto\Delta(\mathcal{A}^{\prime})\}_{c\in\mathcal{C}}$. That is, there exists a $T_{0}$ and $\{\pi_{c}\}_{c\in\mathcal{C}}$ such that the following system of equations holds simultaneously.
$\forall c\in\mathcal{C},\ T^{c}(S^{\prime}|S,a)=\sum_{a^{\prime}}T_{0}(S^{\prime}|S,a,a^{\prime})\pi_{c}(a^{\prime}|S)$ (14)
One can see that the above equation operates independently for each $(S,a)$ pair, but should hold simultaneously for all $c\in\mathcal{C}$ while fixing a pair of $(S,a)$.
For each $(S,a)$, in a matrix notation, Equation (14) can be written as
$M_{C}=M_{\Pi}\cdot M_{T}$
where $M_{C}[c,S^{\prime}]=T^{c}(S^{\prime}|S,a),M_{\Pi}[c,a^{\prime}]=\pi_{c}(a^{\prime}|S),M_{T}[a^{\prime},S^{\prime}]=T_{0}(S^{\prime}|S,a,a^{\prime})$, and therefore, $M_{C}\in\mathbbm{R}^{C\times S},M_{\Pi}\in\mathbbm{R}^{C\times A^{\prime}},M_{T}\in\mathbbm{R}^{A^{\prime}\times S}$. That is, every row $r_{c}\in\mathbbm{R}^{1\times S}$ of $M_{C}$ is a linear combination of all rows in $M_{T}$, and the linear weights is given as the $c$-th row $\in\mathbbm{R}^{1\times A^{\prime}}$ of $M_{\Pi}$. Then, a trivial way to find such $M_{\Pi}$ and $M_{T}$ is as follows,
  1. 1.
    If $|C|\leq|S|$, then one can make $|A^{\prime}|=|C|$, and $M_{\Pi}=\mathbf{I}_{C\times C}$ and $M_{T}=M_{C}$. That is, the policy of the context switcher is to simply select a context in a deterministic way.
  2. 2.
    If $|C|>|S|$, to avoid $M_{\Pi}$ and $M_{T}$ being unboundedly large, a straightforward way is to make $M_{T}=\mathbf{I}_{S\times S}$, i.e. $|A^{\prime}|=|S|$, as the dimension of the row space of $M_{C}$ is at most $|S|$, and therefore, a set of orthonormal basis vectors (i.e., each row of $M_{T}$) can already span the whole row space of $M_{C}$.
One may note that even if in the sub-case when $|C|\leq|S|$, having $|A^{\prime}|=|C|$ may be redundant. A natural question is to find such $M_{\Pi}$ and $M_{T}$ of minimum sizes, i.e. with smallest $|A^{\prime}|$. This further reduces to finding a minimum set of $|S|$-dimensional vectors whose linear combination can represent all the row vectors in $M_{C}$. We now describe a procedure that iteratively construct such $M_{\Pi}$ and $M_{T}$, formally given in Algorithm 1. The idea is quite clean and elegant: start with any arbitrary row in $M_{C}$ as the first basis and an ordering of the rest rows, project the $i$-th subsequent row onto all previous $(i-1)$ basis vectors, and treat the orthogonal residual as the $i$-th basis vector if it is non-zero. One may note that this procedure resembles the Gram-Schmidt Orthogonalization [14], which can be done in strongly-polynomial time. The only difference is, the standard Gram-Schmidt Orthogonalization starts with a set of vectors that are already linearly independent but might not be orthogonal to each other, while here we start with a set of vectors that are possibly linearly dependent and the goal is to find the minimum set of basis vectors.
Algorithm 1 Find the minimum $M_{\Pi}$ and $M_{T}$
1:  Input: $M_{C}$ with $|C|\leq|S|$
2:  Output: $M_{\Pi}$ and $M_{T}$ of the minimum size
3:  Initialize: $M_{T}$ as an empty matrix
4:  $M_{T}.append\_row(M_{C}[1])$
5:  for $i=2$ to $|C|$ do
6:     $new\leftarrow M_{C}[i]-\sum_{k=1}^{i-1}\frac{\langle M_{T}[k],M_{C}[i]\rangle}{\langle M_{T}[k],M_{T}[k]\rangle}M_{T}[k]$
7:     if $new\neq\vec{0}$ then
8:        $M_{T}.append\_row(new)$
9:     end if
10:  end for
11:  $M_{T}\leftarrow normalize\_each\_row(M_{T})$
12:  $M_{\Pi}\leftarrow M_{C}\cdot transpose(M_{T})$
13:  return  $M_{\Pi}$, $M_{T}$
Therefore, one can conclude that the theorem below directly follows from Theorem 7 and Theorem 8.
Theorem 9.
The computational problem of computing the best response to a mixed constant-memory strategy is as hard as that of optimally solving CMDPs.
It has been proved that in infinitely repeated games, there exists a Turing machine strategy such that no Turing machine can implement its best response [27, 35]. Based on the widely accepted belief that CMDP is in the same complexity class as POMDP, Theorem 9 further implies that to render the best response computationally undecidable, it is probably sufficient for the opponent to mix over just a few constant-memory strategies, rather than implementing a complex strategy with perfect memorization via a computer program.
Another remark is, as indicated by Equation (13), if in each turn the opponent is allowed to switch to a different strategy independently of previous actions, then solving such a belief-induced MDP is sufficient for computing the best response, which is definitely decidable and even has constant-memory strategies as optimal policies.
5 Revisiting Single-Agent Decision Processes: A Generative Framework
Building on the theoretical insights from the previous discussion, we now present a generative framework for creating single-agent learning environments. RL is notorious for its poor generalization ability even within the same domain, as an RL model is typically trained for a particular environment simulated by an underlying fixed MDP-like process. Since generalizability is crucial in many real-world applications, especially when an RL algorithm is deployed, the community has investigated some similar research ideas under various names, e.g., multi-task RL [50, 53, 52], contextual RL [23, 11], epistemic POMDPs [20], and zero-shot generalization [26].
Therefore, a framework that can automatically generate environment instances with some shared features is of great significance, since it can be utilized for both training and evaluating the generalizability of a single-agent RL algorithm. There are a few exemplars, e.g. using procedural content generation to make level- or seed-based video games (ProcGen [16]), or simulating existing RL environments with user-specified contextual parameters (CARL [^1] https://github.com/automl/CARL [11]). As a supplement, here we also provide a framework for generating for generating instances of single-agent RL environments. More precisely, this can be done by specifying opponents with controlled memory (hence controlled bounded rationality) under a multi-agent skeleton simulator. The induced decision-making process for the remaining agent will then be a highly configurable learning environment. A key advancement of our generative framework is that the induced single-agent environments can be arbitrarily hard to learn, while the aforementioned existing benchmarks in the literature only diversify tasks with comparable levels of difficulty.
As we addressed from the theoretical side, optimally solving the induced single-agent decision problem involves best responding to the underlying hidden opponents, in other words, the reward/context controller. Once we have an underlying skeleton simulator as a stochastic game (e.g., Figure 1), and specify a context/reward controller as an opponent, we can generate a single-agent learning environment that is arbitrary hard to learn through the following procedures, in ascending order of difficulty:
  1. 1.
    If the opponent plays a single stationary strategy, this will lead to a standard RL environment.
  2. 2.
    If the opponent plays a single behavioral $K$-memory strategy, this will result in an RL environment that requires the learner to at least stack the past $K$ observations, e.g., Figure 2.
  3. 3.
    If the opponent plays a mixed $K$-memory strategy, this will lead to an RL environment that requires the learner to leverage a Bayesian adaptive policy, e.g., Figure 3.
1class StochasticGame:
2 def __init__(self, ...):
3 # instantiate a stochastic game
4 pass
5
6 def reset(self):
7 # initialize a new match, and return the initial state
8 return obs_n, info_n
9
10 def step(self, action_n):
11 # proceed the game to the next state
12 return obs_n, reward_n, term_n, trunc_n, info_n
Figure 1: Standard Pythonic interfaces of a multi-agent simulator as a stochastic game.
1class KMemoryOpponentWrapper:
2 def __init__(self, stochastic_game, agent_i, pi_others, K):
3 # take in a stochastic game as the underlying skeleton
4 self.sg = stochastic_game
5 self.agent_i = agent_i
6 self.pi_others = pi_others
7 self.K = K
8
9 def reset(self):
10 # initialize a new match, and return the initial state
11 obs_n, info_n = self.sg.reset()
12 self.hist_K = [obs_n]
13 return obs_n[self.agent_i], info_n[self.agent_i]
14
15 def step(self, action_i):
16 # proceed the game to the next state
17 action_others = self.pi_others(self.hist_K)
18 obs_n, reward_n, term_n, trunc_n, info_n =\
19 self.sg.step((action_i, action_others))
20 self.hist_K = (self.hist_K + [obs_n])[-self.K:]
21 return (obs_n[self.agent_i], reward_n[self.agent_i],
22 term_n[self.agent_i], trunc_n[self.agent_i],
23 info_n[self.agent_i])
Figure 2: Single-agent learning environments induced by a behavioral $K$-memory opponent strategy.
1class MixedKMemoryOpponentWrapper(KMemoryOpponentWrapper):
2 def __init__(self, stochastic_game, agent_i,
3 pi_others_set, probs, K):
4 # take in a stochastic game as the underlying skeleton
5 super().__init__(stochastic_game, agent_i, None, K)
6
7 # now a set of strategies and the corresponding probs
8 self.pi_others_set = pi_others_set
9 self.probs = probs
10
11 def reset(self):
12 # initialize a new match, and return the initial state
13 obs, info = super().reset()
14 self.pi_others =\
15 random.choice(self.pi_others_set, p=self.probs)
16 return obs, info
Figure 3: Single-agent learning environments induced by a mixed $K$-memory opponent strategy.
6 Conclusion and Future Work
In this work, we comprehensively investigate the concept of $K$-memory strategies. We first establish the best responses and Nash equilibria for behavioral $K$-memory strategies, followed by a discussion on the computational hardness of best responding to mixed $K$-memory strategies. Those theoretic insights later empower a generative framework for studying generalizability of single-agent RL algorithms.
A few research problems still remain open, e.g.,
  1. 1.
    How to efficiently synthesize a truly $K$-memory strategy (ensuring it does not accidentally reduce to a $(K-1)$-memory one), and how to compactly represent it (instead of enumerating all the entries)? Chances are that these two lie in two ends of a spectrum.
  2. 2.
    What can be the minimum effort to make a constant-memory strategy even more powerful? For example, in the Iterated Prisoner’s Dilemma, if both agents adopt a cyclic strategy (constant-memory) together with a Grim-Trigger (definitely not constant-memory but way easy to simulate using a binary indicator), then any payoff profile (under the average-payoff criterion though) can be realized as the outcome of an equilibrium [44].
In summary, we hope this work provide a concise way to characterize bounded rationality in agents that learn and interact, benefiting various real-world applications, such as human-robot cohabitation, strategic LLM-agents, and other domains in cognitive science. We will report more comprehensive experimental results in the near future.